Sunset Sales has 6.8 percent coupon bonds on the market with 11 years left to maturity. The bonds make semiannual payments and currently sell for 98.6 percent of par. What is the effective annual yield?

Multiple Choice

  • 7.24 percent

  • 7.19 percent

  • 7.33 percent

  • 7.11 percent

  • 7.07 percent
Answer #1

For Sunset sales' bond,

Let face valu of bond be 1000

coupon rate = 6.8% compounded semiannually,

So, coupon = 6.8% of 1000/2 = $34

years to maturity = 11 years

current price = 98.6% of 1000 = $986

using financial calculator wirh values

FV = 1000

PV = -986

PMT = 34

N = 2*11 = 22

Solve for I/Y, we get I/Y = 3.49%

So, YTM = 2*3.49 = 6.98% compounded semiannually,

So effective annual rate = (1+APR/2)^2 - 1 = (1 + 0.0698/2)^2 - 1 = 7.11%

So, option D is correct.